# Investigating the role of mutual information in the Page curve for a functional renormalization group improved Schwarzschild black hole

### v1

@@ Abstract
The present work delves into probing the importance of mutual information of relevant subsystems in obtaining the correct time-evolution of fine-grained entropy of Hawking radiation, as suggested by Page. This was done by considering a functional renormalization group improved or simply, quantum corrected Schwarzschild black solution which captures the flavour of an effective theory of quantum gravity. The mentioned black hole solution emerges from a asymptotically safe average effective action which describes a trajectory in momentum-space and satisfies a renormalizarion group equation. Furthermore, in the before Page time scenario, the behaviour of the mutual information between appropiate subsystems over time leads to the Hartman-Maldacena time. The observations made for the quantum corrected Schwarzschild black hole have been compared to the same made for the standard Schwarzschild black hole in order to draw some novel conclusions. Based upon our observations, we also propose a formula for computing fine-grained entropy of Hawking radiation for an eternal black hole, in presence of the island.
@@
However, as the origin of the Hawking radiation relies upon the fundamental concepts of quantum physics, it will be a matter basic instinct to probe the relevant measurements of quantum statistical mechanics also. Historically, this quest has Hawking to the computation of von-Neumann entropy Nielsen and Chuang (2000) of the black hole radiation which is a fine-grained notion of entropy unlike the usual coarse-grained notion of entropy (Bekenstein-Hawking entropy). In Hawking (1976), he showed that the von-Neumann entropy of Hawking radiation is a monotonically increasing function with respect to time of an asymptotic observer but from the landscape of quantum mechanical postulates, this statement creates a paradoxical scenario. This can be understood in the following way. It is known that the formation of black holes are (due to the collapse of a massive shell) corresponds to a pure state which is represented by the fact that at the beginning of the evaporation process, the von-Neumann entropy of radiation is zero. Moreover, at the end of the evaporation process, the von-Neumann entropy of radiation once again should be zero. This is quite a expected result from the perspective of unitary evolution of the quantum state over time. This contradictory situation can be denoted as the entropic version of the information paradox and the best way to understand the scenario is to take help of the Page curve. The Page curve describes the time evolution of the von Neumann entropy of Hawking radiation Page (1993). One can also naively say that this paradoxical situation can be solved if one can obtain the Page curve of radiation which abide by the rules of quantum mechanics. Before we indulge in to the subtleties of the Page curve it should be mentioned that, the von Neumann entropy of the Hawking radiation is identified with that of the matter fields residing on the outside region of the black hole, namely, region $R$. Now, according to the Page curve, starting from zero, the von-Neumann entropy of the radiation $(S_{vN}(R))$ increases over time. On the other hand, due to the evaporation the thermal entropy of black hole decreases with time and at a particular time (the Page time $t_{P}$) it becomes equal to the thermal entropy of the black hole (that is the Bekenstein-Hawking entropy $S_{BH}$). This can also be depicted as fine-grained entropy achieving the coarse-grained or maximal value. It can be noted that upto Page time there is no paradoxical situation because in this time domain $(t\leq t_{P})$, the fine-grained entropy of radiation is bounded by the coarse-grained (thermal) entropy of the black hole, that is $S_{vN}(R)\leq S_{BH}$. However, after the Page time, for $t>t_{P}$ the von-Neumann entropy of Hawking radiation is greater than the thermal entropy of the black hole. Thus the paradox appears just after the Page time and therefore the challenge is to obtain the correct behaviour of fine-grained entropy after the Page time. As we have mentioned earlier, the unitarity principle suggests that $S_{vN}(R)$ should decrease after the Page time and at the end of the evaporation it should vanish Page (1993, 2013). Some interesting observations in this directions can be found in Almheiri et al. (2013a, b); Lloyd and Preskill (2014); Papadodimas and Raju (2014).
@@
Recent advancements of the AdS/CFT correspondence has provided us a fresh insight to the above discussed scenario. It has been suggested that while computing the von-Neumann entropy of radiation, we have to take into account the contribution of certain region formed inside black hole horizon along with the traditional contributions coming from the matter fields located outside the black hole, where the radiation is collected. This region which is formed in the interior of the black hole is known as the islands Penington et al. (2022); Penington (2020); Almheiri et al. (2020a, 2019a, b, 2019b, 2021), and the end points of the island surface is called the quantum extremal surfaces (QES) Engelhardt and Wall (2015); Engelhardt and Fischetti (2019); Akers et al. (2020); Wall (2014). Keeping in mind this additional contribution, the fine-grained entropy of Hawking radiation in the presence of island is given by
@@
In a simple sense, the first term in the above formula represents the geometrical contributions and the second term summarizes the contribution coming from the von-Neumann entropy of matter fields localized on the region $R\cup I$. The main essence of the above formula lies in the fact that it provides us a non-trivial generalization of the fine-grained entropy $S_{vN}(R)$. It can be observed that in order to obtain the correct value of $S_{vN}(R)$, one has to perform extremization over all possible island configurations $I$ and has to the consider the minimum of all extremal values, if there is more than one extrema.
@@
The island formula has emerged from the application of replica trick to compute the von-Neumann entropy of matter fields, in presence of a dynamical gravitational background. In this method, the partition function appears as a gravitational path integral. Apart from the previously known Hawking saddle-point, it has been noted that there exists an additional non-trivial saddle-point known as the replica wormhole saddle-point [^1] A detailed review regarding wormholes and their utility in gauge/gravity duality can be found in Kundu (2022). (with proper boundary conditions) which acts as the dominating saddle-point depending under certain condition Almheiri et al. (2020b); Penington et al. (2022); Almheiri et al. (2020b); Goto et al. (2021); Colin-Ellerin et al. (2021). This new formula of computing the fine-grained of Hawking radiation produces the Page curve which one expects from the principle of unitarity and due to this remarkable observation, the island formulation has emerged as an important prescription to be studied in recent times Chen et al. (2020); Hashimoto et al. (2020); Hartman et al. (2020); Anegawa and Iizuka (2020); Dong et al. (2020); Balasubramanian et al. (2021); Alishahiha et al. (2021); Azarnia et al. (2021); Aref’eva and Volovich (2021); He et al. (2022); Omidi (2022); Yu et al. (2022); Yadav (2022); Du et al. (2022). However, there are certain issues regarding this approach which needs to be properly addressed Geng and Karch (2020); Geng et al. (2022a).
@@
This formalism was first applied to the black hole solution of the JT gravity filled with 2d conformal matter Almheiri et al. (2020a); Hollowood and Kumar (2020); Goto et al. (2021). In early time scenario, that is, $t\leq t_{P}$, the Hawking saddle-point of the gravitational path integral dominates and the fine-grained entropy of Hawking radiation increases with respect to time. At the Page time $t=t_{P}$, it becomes equal to the value of coarse-grained entropy of radiation. However, just after the Page time, that is for $t>t_{P}$, the replica wormhole saddle-point starts to dominate and the contribution from the interior region ($I$) becomes significant which eventually lead us to the expected Page curve. There are many interesting developments in recent years in this directions, such as 2d asymptotically flat or AdS black holes Anegawa and Iizuka (2020); Hartman et al. (2020); Wang et al. (2021a); Li et al. (2021); Pedraza et al. (2021), higher dimensional black holes Almheiri et al. (2020c); He et al. (2022); Alishahiha et al. (2021); Hashimoto et al. (2020); Matsuo (2021); Krishnan (2021); Krishnan et al. (2020); Omidi (2022); Geng et al. (2021a, 2022b); Goswami and Narayan (2022); Du et al. (2022); Chu et al. (2021), charged black holes Ling et al. (2021); Wang et al. (2021b); Kim and Nam (2021); Ahn et al. (2022); Yu and Ge (2022) etc. Additional interesting studies can be found in Balasubramanian et al. (2021); Chen et al. (2021); Azarnia et al. (2021); Geng et al. (2021b); Bhattacharya et al. (2021); Yu and Ge (2024); Gomez (2024); Jain et al. (2023); Guo and Miao (2023); Nikolakopoulou ; Anand (2023a); Chang et al. (2023); Li and Miao (2023); Guo et al. (2023); Miao (2023); Yadav and Joshi (2023); Hosseini Mansoori et al. (2022); Hu et al. (2022); Yu and Ge (2023); Goswami and Narayan (2022); Djordjević et al. (2022); Anand (2023b); Azarnia and Fareghbal (2022); Seo (2022); Yu and Ge (2024); Wu et al. (2024); Yu et al. (2023); Luongo et al. (2023).
@@ I Asymptotically safe theory of quantum gravity and the quantum corrected Schwarzschild black hole
$ds^{2}=-\left(1-\frac{2GM}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2GM}{r}\right)}+r^{2}d\Omega_{2}^{2}$ (4)
where $G$ is the Newton’s gravitational constant and $M$ represents the mass of the black hole. The Schwarzschild solution is very useful to describe many physical phenomena. This solution of Einstein equation is purely classical. The horizon radius of the Schwarzschild black hole is given by $r^{(S)}=2GM$. The Hawking temperature $(T_{H}^{(S)})$ and the entropy $(S_{BH}^{(S)})$ of the Schwarzschild black hole are given by
$T_{H}^{(S)}$ $=$ $\frac{1}{4\pi r^{(S)}}=\frac{1}{8\pi GM}$ (5)
$S_{BH}^{(S)}$ $=$ $\frac{4\pi(r^{(S)})^{2}}{4G}$ (6)
@@
where $r_{+}$ and $r_{-}$ represent the outer and inner horizons respectively. We can rewrite the lapse function in terms of $r_{+}$ and $r_{-}$ in the following way
$f(r)=\frac{1}{r^{2}}(r-r_{+})(r-r_{-})\leavevmode\nobreak\ .$ (13)
@@
In this work we are only interested in the black hole sector. For the subsequent analysis, we will use the above form of horizon radius. The surface gravity and then Hawking temperature of QCSBH  [^3] The inverse of $T_{H}^{(q)}$, $\beta^{(q)}=8\pi GM\left(1+\frac{\tilde{\omega}^{2}}{4M^{4}G^{2}}\right)=\beta^{(S)}\left(1+\frac{\tilde{\omega}^{2}}{4M^{4}G^{2}}\right)$, where $\beta^{(S)}=8\pi GM,$ is the inverse temperature of the classical Schwarzschild black hole.is given by
$\kappa^{(q)}$ $=$ $\frac{1}{4G_{0}M}\left(1-\frac{\tilde{\omega}^{2}}{4M^{4}G_{0}^{2}}\right)$ (16)
$T_{H}^{(q)}$ $=$ $T_{H}^{(S)}\left(1-\frac{\tilde{\omega}^{2}}{4M^{4}G_{0}^{2}}\right)$ (17)
@@
$r^{*}(r)$ $=$ $\int\frac{dr}{f(r)}$
$=$ $r+\frac{r_{+}^{2}}{r_{+}-r_{-}}\log(r-r_{+})-\frac{r_{-}^{2}}{r_{+}-r_{-}}\log(r-r_{-})$
@@
By using the metric in Kruskal coordinate, we can proceed to construct the Penrose diagram for a two-sided eternal QCSBH geometry. This we provide in Fig.(1).
@@
We now proceed to explore the proposals given in Saha et al. (2022); Roy Chowdhury et al. (2022, 2023) for the black hole solution discussed above.

### v2

@@ Abstract
The present work delves into probing the importance of mutual information of relevant subsystems in obtaining the correct time-evolution of fine-grained entropy of Hawking radiation, as suggested by Page. This was done by considering a functional renormalization group improved or simply, quantum corrected Schwarzschild black solution which captures the flavour of an effective theory of quantum gravity. The mentioned black hole solution emerges from an asymptotically safe average effective action which describes a trajectory in momentum-space and satisfies a renormalizarion group equation. Furthermore, in the before Page time scenario, the behaviour of the mutual information between appropiate subsystems over time leads to the Hartman-Maldacena time. The observations made for the quantum corrected Schwarzschild black hole have been compared to the same made for the standard Schwarzschild black hole in order to draw some novel conclusions. Based upon our observations, we also propose a formula for computing fine-grained entropy of Hawking radiation for an eternal black hole, in presence of the island.
@@
However, as the origin of the Hawking radiation relies upon the fundamental concepts of quantum physics, it will be a matter of basic instinct to probe the relevant measurements of quantum statistical mechanics also. Historically, this quest has Hawking to the computation of von-Neumann entropy Nielsen and Chuang (2000) of the black hole radiation which is a fine-grained notion of entropy unlike the usual coarse-grained notion of entropy (Bekenstein-Hawking entropy). In Hawking (1976), he showed that the von Neumann entropy of Hawking radiation is a monotonically increasing function with respect to time of an asymptotic observer but from the landscape of quantum mechanical postulates, this statement creates a paradoxical scenario. This can be understood in the following way. It is known that the formation of black holes are (due to the collapse of a massive shell) corresponds to a pure state which is represented by the fact that at the beginning of the evaporation process, the von Neumann entropy of radiation is zero. Moreover, at the end of the evaporation process, the von Neumann entropy of radiation once again should be zero. This is quite a expected result from the perspective of unitary evolution of the quantum state over time. This contradictory situation can be denoted as the entropic version of the information paradox and the best way to understand the scenario is to take help of the Page curve. The Page curve describes the time evolution of the von Neumann entropy of Hawking radiation Page (1993). One can also naively say that this paradoxical situation can be solved if one can obtain the Page curve of radiation which abide by the rules of quantum mechanics. Before we indulge in to the subtleties of the Page curve it should be mentioned that, the von Neumann entropy of the Hawking radiation is identified with that of the matter fields residing on the outside region of the black hole, namely, region $R$. Now, according to the Page curve, starting from zero, the von-Neumann entropy of the radiation $(S_{vN}(R))$ increases over time. On the other hand, due to the evaporation the thermal entropy of black hole decreases with time and at a particular time (the Page time $t_{P}$) it becomes equal to the thermal entropy of the black hole (that is the Bekenstein-Hawking entropy $S_{BH}$). This can also be depicted as fine-grained entropy achieving the coarse-grained or maximal value. It can be noted that upto Page time there is no paradoxical situation because in this time domain $(t\leq t_{P})$, the fine-grained entropy of radiation is bounded by the coarse-grained (thermal) entropy of the black hole, that is $S_{vN}(R)\leq S_{BH}$. However, after the Page time, for $t>t_{P}$ the von Neumann entropy of Hawking radiation is greater than the thermal entropy of the black hole. Thus the paradox appears just after the Page time and therefore the challenge is to obtain the correct behaviour of fine-grained entropy after the Page time. As we have mentioned earlier, the unitary principle suggests that $S_{vN}(R)$ should decrease after the Page time and at the end of the evaporation it should vanish Page (1993, 2013). Some interesting observations in this directions can be found in Almheiri et al. (2013a, b); Lloyd and Preskill (2014); Papadodimas and Raju (2014).
@@
Recent advancements of the AdS/CFT correspondence has provided us a fresh insight to the above discussed scenario. It has been suggested that while computing the von Neumann entropy of radiation, we have to take into account the contribution of certain region formed inside black hole horizon along with the traditional contributions coming from the matter fields located outside the black hole, where the radiation is collected. This region which is formed in the interior of the black hole is known as the islands Penington et al. (2022); Penington (2020); Almheiri et al. (2020a, 2019a, b, 2019b, 2021), and the end points of the island surface is called the quantum extremal surfaces (QES) Engelhardt and Wall (2015); Engelhardt and Fischetti (2019); Akers et al. (2020); Wall (2014). Keeping this additional contribution in mind, the fine-grained entropy of Hawking radiation in the presence of island is given by
@@
In a simple sense, the first term in the above formula represents the geometrical contributions and the second term summarizes the contribution coming from the von Neumann entropy of matter fields localized on the region $R\cup I$. The main essence of the above formula lies in the fact that it provides us a non-trivial generalization of the fine-grained entropy $S_{vN}(R)$. It can be observed that in order to obtain the correct value of $S_{vN}(R)$, one has to perform extremization over all possible island configurations $I$ and has to the consider the minimum of all extremal values, if there is more than one extrema.
@@
The island formula has emerged from the application of replica trick to compute the von-Neumann entropy of matter fields, in presence of a dynamical gravitational background. In this method, the partition function appears as a gravitational path integral. Apart from the previously known Hawking saddle-point, it has been noted that there exists an additional non-trivial saddle-point known as the replica wormhole saddle-point [^1] A detailed review regarding wormholes and their utility in gauge/gravity duality can be found in Kundu (2022). (with proper boundary conditions) which acts as the dominating saddle-point depending under certain condition Almheiri et al. (2020b); Penington et al. (2022); Almheiri et al. (2020b); Goto et al. (2021); Colin-Ellerin et al. (2021). This new formula of computing the fine-grained of Hawking radiation produces the Page curve which one expects from the principle of unitarity and due to this remarkable observation, the island formulation has emerged as an important prescription to be studied in recent times Chen et al. (2020); Hashimoto et al. (2020); Hartman et al. (2020); Anegawa and Iizuka (2020); Dong et al. (2020); Balasubramanian et al. (2021); Alishahiha et al. (2021a); Azarnia et al. (2021); Aref’eva and Volovich (2021); He et al. (2022); Omidi (2022); Yu et al. (2022); Yadav (2022); Du et al. (2022). However, there are certain issues regarding this approach which needs to be properly addressed Geng and Karch (2020); Geng et al. (2022a).
@@
This formalism was first applied to the black hole solution of the JT gravity filled with 2d conformal matter Almheiri et al. (2020a); Hollowood and Kumar (2020); Goto et al. (2021). In early time scenario, that is, $t\leq t_{P}$, the Hawking saddle-point of the gravitational path integral dominates and the fine-grained entropy of Hawking radiation increases with respect to time. At the Page time $t=t_{P}$, it becomes equal to the value of coarse-grained entropy of radiation. However, just after the Page time, that is for $t>t_{P}$, the replica wormhole saddle-point starts to dominate and the contribution from the interior region ($I$) becomes significant which eventually lead us to the expected Page curve. There are many interesting developments in recent years in this directions, such as 2d asymptotically flat or AdS black holes Anegawa and Iizuka (2020); Hartman et al. (2020); Wang et al. (2021a); Li et al. (2021); Pedraza et al. (2021), higher dimensional black holes Almheiri et al. (2020c); He et al. (2022); Alishahiha et al. (2021a); Hashimoto et al. (2020); Matsuo (2021); Krishnan (2021); Krishnan et al. (2020); Omidi (2022); Geng et al. (2021a, 2022b); Goswami and Narayan (2022); Du et al. (2022); Chu et al. (2021), charged black holes Ling et al. (2021); Wang et al. (2021b); Kim and Nam (2021); Ahn et al. (2022); Yu and Ge (2022) etc. Some other interesting studies in this direction can also be found in Balasubramanian et al. (2021); Chen et al. (2021); Azarnia et al. (2021); Geng et al. (2021b); Bhattacharya et al. (2021); Yu and Ge (2024); Gomez (2024); Jain et al. (2023); Guo and Miao (2023); Nikolakopoulou ; Anand (2023a); Chang et al. (2023); Li and Miao (2023); Guo et al. (2023); Miao (2023); Yadav and Joshi (2023); Hosseini Mansoori et al. (2022); Hu et al. (2022); Yu and Ge (2023); Goswami and Narayan (2022); Djordjević et al. (2022); Anand (2023b); Azarnia and Fareghbal (2022); Seo (2022); Yu and Ge (2024); Wu et al. (2024); Yu et al. (2023); Luongo et al. (2023); Uhlemann (2021); Karch et al. (2022).
@@ I Asymptotically safe theory of quantum gravity and the quantum corrected Schwarzschild black hole
$ds^{2}=-\left(1-\frac{2G_{0}M}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2G_{0}M}{r}\right)}+r^{2}d\Omega_{2}^{2}$ (4)
where $G_{0}$ is the Newton’s gravitational constant and $M$ represents the mass of the black hole. The Schwarzschild solution is very useful to describe many physical phenomena. This solution of Einstein equation is purely classical. The horizon radius of the Schwarzschild black hole is given by $r^{(S)}=2G_{0}M$. The Hawking temperature $(T_{H}^{(S)})$ and the entropy $(S_{BH}^{(S)})$ of the Schwarzschild black hole are given by
$T_{H}^{(S)}$ $=$ $\frac{1}{4\pi r^{(S)}}=\frac{1}{8\pi G_{0}M}$ (5)
$S_{BH}^{(S)}$ $=$ $\frac{4\pi(r^{(S)})^{2}}{4G_{0}}$ (6)
@@
where $r_{+}$ and $r_{-}$ represent the outer and inner horizons respectively. It is also important to note that the number of horizons depends sensitively on the mass parameter. In fact, it can be shown that there exists a critical mass $M_{\text{cr}}$ that governs the horizon structure, for masses greater than $M_{\text{cr}}$, the black hole exhibits two distinct horizons, while for masses below this threshold, no horizon exists, and if $M=M_{cr}$ there is only one horizon. In this work, we have considered the scenario where $M>M_{\text{cr}}$, which implies the existence of two distinct horizons Koch and Saueressig (2014). Therefore, we can rewrite the lapse function in terms of $r_{+}$ and $r_{-}$ in the following way
$f(r)=\frac{(r-r_{+})(r-r_{-})}{r^{2}+\tilde{\omega}G_{0}}\leavevmode\nobreak\ .$ (13)
@@
In this work we are only interested in the black hole sector. For the subsequent analysis, we will use the above form of horizon radius. The surface gravity and then Hawking temperature of QCSBH  [^3] The inverse of $T_{H}^{(q)}$, $\beta^{(q)}=8\pi G_{0}M\left(1+\frac{\tilde{\omega}G_{0}}{(r^{(s)})^{2}}\right)=\beta^{(S)}\left(1+\frac{\tilde{\omega}G_{0}}{(r^{(s)})^{2}}\right)$, where $\beta^{(S)}=8\pi GM,$ is the inverse temperature of the classical Schwarzschild black hole.is given by
$\kappa^{(q)}$ $=$ $\frac{1}{4G_{0}M}\left(1-\frac{\tilde{\omega}G_{0}}{(r^{(s)})^{2}}\right)$ (16)
$T_{H}^{(q)}$ $=$ $T_{H}^{(S)}\left(1-\frac{\tilde{\omega}G_{0}}{(r^{(s)})^{2}}\right)$ (17)
@@
$r^{*}(r)$ $=$ $\int\frac{dr}{f(r)}$ (20)
$=$ $r+\left(\frac{r_{+}^{2}+\tilde{\omega}G_{0}}{r_{+}-r_{-}}\right)\log(r-r_{+})$
$-\left(\frac{r_{-}^{2}+\tilde{\omega}G_{0}}{r_{+}-r_{-}}\right)\log(r-r_{-})\leavevmode\nobreak\ .$
@@
By using the metric in Kruskal coordinate, we can proceed to construct the Penrose diagram for a two-sided eternal QCSBH geometry. This we provide in Fig.(1). The horizon structure of the quantum-corrected Schwarzschild black hole is indeed very similar to that of the Reissner-Nordström (RN) black hole. However, it is important to note that while the RN black hole arises from adding a gauge field to the Einstein-Hilbert action, the black hole we have considered is obtained from the functional renormalization group approach. Further, expanding eq.(13) upto $\mathcal{O}(\tilde{\omega})$, the lapse function has the form
$f(r)\approx 1-\frac{2MG_{0}}{r}+\frac{2M\tilde{\omega}G_{0}^{2}}{r^{3}}\leavevmode\nobreak\ .$ (23)
This is significantly different from the RN black hole. Hence, the physics for the QCSBH and the RN black hole are quite different.
We now proceed to explore the proposals given in Saha et al. (2022); Roy Chowdhury et al. (2022, 2023) for the black hole solution discussed above. It is to be noted that qualitatively the QCSBH solution is an improved solution resulting from the flow of the Newton’s gravitational constant $G$ arising from the functional renormalization group approach to quantum gravity. The flow equation for $G$ following from this approach does not introduce any higher curvature corrections to the Einstein-Hilbert action which can potentially modify the area term of the island formula (given in eq.(1)), as shown in Alishahiha et al. (2021b). Some previous studies related to island formalism for modified theories of gravity can be found in Tong et al. (2024); Anand (2023a); Jain et al. (2023); Omidi (2022); Yadav (2022); Anand (2023b); Yadav and Misra (2023); Djordjević et al. (2022); Liu et al. (2025); Dong and Tian (2025); Alishahiha et al. (2021a); Lu and Lin (2022).